A particle moves so that its position vector is given by r =cosωti+sinωtj. where ω is a constant.


a)Show that the velocity v of the particle is perpendicular to r, and the velocity acceleration a is directed towards the origin and has magnitude proportional to the distance from the origin.


b) Show also that r × r is a constant vector.


1
Expert's answer
2021-06-10T02:50:52-0400

Given "\overrightarrow{r}= cos\omega \hat{x} + sin\omega\hat{y} ........(1)"

Velocity, "\overrightarrow{v} = \dfrac{d\overrightarrow{r}}{dt}= - \omega sin\omega t\hat{x} + \omega cos\omega\hat{y}.........(2)"

Acceleration, "\overrightarrow{a} = \dfrac{d^2\overrightarrow{r}}{dt^2} = -\omega^2 cos\omega t \hat{x} - \omega^2sin\omega t \hat{y} = -\omega^2\overrightarrow{r} ............(3)"

Now using (1),(2) and (3) we get, "\overrightarrow{v}.\overrightarrow{r} =0" Its indicate that velocity is perpendicular to "\overrightarrow{r}" and "\overrightarrow{a}.\overrightarrow{r}\neq 0" So acceleration is neither parallel nor perpendicular to "\overrightarrow{r}"

The equation (3) indicates the acceleration is directed toward the origin.


(b) Ans:-

"\overrightarrow{r}= cos\omega \hat{x} + sin\omega\hat{y}"


"|\overrightarrow{r}|=\sqrt{(Cos\omega)^2+(Sin\omega)^2}=1"


As we know that the angle between itself will be "0\degree"

Hence,

"\overrightarrow{r}\times \overrightarrow{r}= |\overrightarrow{r}| .|\overrightarrow{r}|.Sin0\degree=0"


Hence, "\overrightarrow{r}\times \overrightarrow{r}" is a constant vector.